# Invariant Features for Global Crop Type Classification

### v1

@@ 2.2 Reference Data
Australia (AUS) — Queensland Seasonal Crop (QSC): QSC [42] is a seasonal crop classification product used to monitor large-scale cropping systems in Queensland, with the original data provided in vector format. Since 1988, multi-spectral time-series imagery has been used to generate two seasonal maps per year—one for winter crops (June–October) and another for summer crops (November–May). The 2023 summer crop map used in this study includes cotton, sugarcane, and an “Other” class that may contain corn, soybeans, and additional crop types. Reported accuracies are high, with user accuracy around 95% and producer accuracy about 90% for major crops [43]. Data access: https://www.qld.gov.au/environment/land/management/mapping/statewide-monitoring/crops.
@@ 2.3 Sample Collection
For hyperspectral data, we used the EMIT Level-2A Estimated Surface Reflectance product (available via the GEE collection ‘NASA/EMIT/L2A/RFL’), which contains estimated surface reflectance, associated uncertainty, and cloud mask information. The built-in buffered cloud mask was used to exclude all pixels with cloud mask value equal to 1. To evaluate the utility of hyperspectral observations at different time points, we selected representative months that align with key crop growth stages, based on EMIT data availability over agricultural regions. For Northern Hemisphere countries, we used imagery from June and August 2023. We examined EMIT acquisitions throughout the main crop growing season and found that the overlap of cloud-free observations was largest between June and August. Including additional months such as May, July, or September resulted in negligible overlap and thus provided virtually no usable data. Median reflectance values were calculated for each month, and a combined availability mask—defined as the intersection of the two monthly cloud-free masks—was applied to retain valid pixels. For Southern Hemisphere countries, the timeframe was shifted by six months. As EMIT only established comprehensive spatial coverage in 2023, no November 2022 data were available over our study regions, and thus only February 2023 was used. In regions with data availability, monthly median hyperspectral reflectance was calculated using GEE.
@@
For ARG, AUS, and CHN, the original datasets are binary or three-class products, and thus no further class selection or merging was required.
For datasets originally provided in vector format, namely QSC (AUS), LGP (BEL), RPG (FRA), CROME (GBR), and BRP (NLD), we converted them into raster format at a spatial resolution of 10 meters and reprojected them to the WGS84 coordinate reference system. To further improve label reliability, we applied a 60-meter inward buffer (based on EMIT resolution) to all labeled parcels after importing the crop maps into GEE, which helped eliminate small noisy patches. As shown in Fig. 2 (a), erroneous labels distributed within sample areas are removed after applying buffer filtering.
@@
Grid-Based Sampling: We matched the buffered crop maps with the filtered satellite data (processed separately for S2 and EMIT) in both space and time. The intersection of valid crop labels and available cloud-free pixels defined the final sampling region, ensuring each pixel had both a class label and satellite data. To reduce sample correlation, the study region was divided into $0.1^{\circ}\times 0.1^{\circ}$ cells in geographic coordinates (WGS84), and random samples were drawn within each cell at a spatial resolution of 60 meters (i.e., ensuring a minimum distance of 60 meters between any two sampled points), as shown in Fig. 2 (b). In USA, one sample was drawn per cell, while in other countries with smaller study areas, five samples were drawn per cell. Taking FRA as an example, a $0.1^{\circ}\times 0.1^{\circ}$ cell is approximately 7.7 km $\times$ 11.1 km, and five uniformly random samples with a minimum of 60-meter spacing remain spatially distinct.
@@ 3 Methodology
To address these challenges, we design a lightweight yet effective CNN model named CropNet, tailored for learning robust features from pixel-level data. Furthermore, to improve the invariance of learned features across geographical domains, where both spectral responses and phenological phases may vary, we introduce a dedicated data augmentation strategy for multi-spectral time-series, aimed at enhancing robustness to temporal and spectral perturbations.
@@ 3.1 Construction of Temporal Multi-Spectral Data
1D/2D median features: To construct consistent median features from raw S2 observations, we apply a windowed median resampling strategy. For each pixel, we define a time span of interest with a start day-of-year (DOY) $T_{s}$ and an end DOY $T_{e}$, which typically covers the crop growing season. Although data are collected throughout the year, not all periods contribute equally to crop type classification. The selected time span allows the model to focus on the most relevant phenological phases.
This time span is divided into equal-length temporal windows of size $d$, resulting in $t=(T_{e}-T_{s})/d$ discrete bins. For each bin, if valid (non-cloudy) observations are available, we compute the median reflectance across all dates within the bin. If no valid observations are found due to cloud coverage, the missing value is filled using linear interpolation along the temporal axis. If the time series contains too many gaps to allow reliable interpolation, the sample is discarded.
Each pixel is ultimately represented as a one-dimensional feature vector of length $10\times t$, where 10 corresponds to the selected S2 bands, and $t$ is the number of time bins.
@@ 3.2 Augmentation of Temporal Multi-Spectral Data
Time Shift simulates phenological shifts by randomly moving the entire time window forward or backward by 0 to 10 days (see Section 4.5 for details). Given the original day-of-year (DOY) range $T_{s},T_{e}]$, a random offset is sampled and applied to both start and end DOYs, resulting in a shifted range $T_{s}^{\prime},T_{e}^{\prime}]$. Median values are then recalculated over this new interval using the same number of time bins as the original, ensuring feature length consistency. This augmentation accounts for planting date differences or climatic variations across regions.
Time Scale adjusts the temporal extent of the sampling window to mimic variability in crop development duration. The start and end DOYs are independently perturbed by values sampled from the range [30,+10][-30,+10] (see Section 4.5 for details), potentially expanding or contracting the time span by up to 40 days. To maintain a fixed input dimension, the temporal window is recomputed such that the total number of bins remains unchanged. This method captures differences in growing season length due to crop varieties, management practices or climate.
Magnitude Warping perturbs the shape of each spectral band’s temporal curve by applying a smooth, multiplicative modulation, aiming to mimic real-world deviations such as sensor noise, sub-pixel heterogeneity, or biomass variability (e.g., the same crop may even have different biomass in different fields) [52]. For median features of each sample, a random warping function is generated by cubic spline interpolation over a few anchor points (knots), whose values are drawn from a Gaussian distribution centered at 1.0. This function is then applied to scale the temporal profile of each spectral band. Based on empirical analysis (Section 4.5), we set the standard deviation of the Gaussian distribution to 0.2 and the number of anchor points (knots) to 5. To avoid over-regularization, each sample undergoes the same warping transformation across all spectral bands. This augmentation helps the model learn more invariant and generalizable representations under varying growing conditions.
@@ 3.3 Classification Model: CropNet
We design a lightweight CNN, named CropNet, specifically for pixel-level classification of crop type. It performs convolutions jointly along the spectral and temporal dimensions to capture dependencies across both domains, as presented in Fig. 6. CropNet consists of four convolutional blocks, each containing two $3\times 3$ convolutional layers with batch normalization and ReLU (Rectified Linear Unit) activations. Down sampling is performed in the first and third blocks via stride-2 convolutions, retaining more temporal resolution while still allowing the model to capture multi-scale spectral-temporal patterns. After the final convolutional stage, global average pooling reduces the spatial dimensions to $1\times 1$, followed by a linear layer for classification. The output is a probability distribution over crop types.
@@
To improve regularization and prevent overfitting on small per-pixel features, we apply spatial dropout [53] after each convolutional block. Unlike standard dropout, which randomly drops individual values, spatial dropout zeros out entire feature channels, forcing the network to learn more robust and distributed representations across the spectrum-time space. This is particularly useful in our case, as the input tensor lacks contextual spatial structure and is more prone to overfitting due to limited feature diversity per sample.
We evaluated several architectural variants with different depths and channel widths. Our experiments showed that a four-stage, nine-layer configuration achieves the best trade-off between performance and model complexity. It provides stable accuracy across test regions and is a minimal design for optimal generalization. Deeper networks showed no significant gains and were more prone to overfitting, while shallower ones consistently underperformed.
For one-dimensional inputs or datasets structured as flat time-series vectors, such as harmonic coefficients or hyperspectral reflectance, the 2D convolutions in CropNet can be easily replaced with 1D convolutions. In this configuration, the network architecture remains structurally identical, treating the input as a single-channel signal and performing convolutions along the feature axis.
@@ 4.1 Experimental Setup
Class alignment: Since available crop categories differ across countries and the definition of the “Other” class also varies, we apply a label alignment strategy for fair evaluation. For AUS, ARG, and CHN [42, 41, 45], where the datasets are binary or ternary and the “Other” class may implicitly include target classes (e.g., soybeans and wheat may exist in AUS “Other”), we adopt the following rule: when using these datasets as test regions, any predicted class not explicitly defined within their available categories is reassigned to “Other” before computing accuracy. For example, samples predicted as soybean or wheat in AUS are reassigned to “Other” during evaluation. In the FRA → USA experiment, since FRA contains no sugarcane or cotton samples, we remove these two classes from USA when conducting evaluation; in this case, they cannot simply be merged into “Other” because FRA entirely lacks these categories. For all other transfer scenarios, no special category alignment is applied, as the training regions cover all classes present in the test regions and the definition of “Other” remains consistent.
@@
To evaluate the intrinsic discriminative ability of different features and models without the added complexity of domain shifts, we conduct in-region classification experiments where training and testing samples are drawn from the same country. Table 4 summarizes the results where models are trained and tested within the same country. For each country, 80% of samples are randomly selected for training and 20% for testing, with experiments repeated ten times. The results show that median features consistently outperform harmonic coefficients, and 2D median features further outperform 1D median features. Data augmentation remains effective even under these in-region settings, improving both model stability and accuracy, despite limited domain shifts between training and testing data.
@@
It is noteworthy that the FRA dataset is based on farmer-reported data and thus highly reliable, while the USA dataset is automatically generated and its label accuracy is inherently imperfect. Despite this, our model achieves near 99% accuracy on the USA data, suggesting that some degree of overfitting to label noise may have occurred. However, the strong generalization observed in USA-to-other transfer experiments suggests that the learned features still generalize well despite noisy labels.

### v2

@@ 2.2 Reference Data
Australia (AUS) — Queensland Seasonal Crop (QSC): QSC [42] is a seasonal crop classification product used to monitor large-scale cropping systems in Queensland, with the original data provided in vector format. Since 1988, multi-spectral time-series imagery has been used to generate two seasonal maps per year, one for winter crops (June–October) and another for summer crops (November–May). The 2023 summer crop map used in this study includes cotton, sugarcane, and an “Other” class that may contain corn, soybeans, and additional crop types. Reported accuracies are high, with user accuracy around 95% and producer accuracy about 90% for major crops [43]. Data access: https://www.qld.gov.au/environment/land/management/mapping/statewide-monitoring/crops.
@@ 2.3 Sample Collection
For hyperspectral data, we used the EMIT Level-2A Estimated Surface Reflectance product (available via the GEE collection ‘NASA/EMIT/L2A/RFL’), which contains estimated surface reflectance, associated uncertainty, and cloud mask information. The built-in buffered cloud mask was used to exclude all pixels with cloud mask value equal to 1. To evaluate the utility of hyperspectral observations at different time points, we selected representative months that align with key crop growth stages, based on EMIT data availability over agricultural regions. For Northern Hemisphere countries, we used imagery from June and August 2023. We examined EMIT acquisitions throughout the main crop growing season and found that the overlap of cloud-free observations was largest between June and August. Including additional months such as May, July, or September resulted in negligible overlap and thus provided virtually no usable data. Median reflectance values were calculated for each month, and a combined availability mask, which was defined as the intersection of the two monthly cloud-free masks, was applied to retain valid pixels. For Southern Hemisphere countries, the timeframe was shifted by six months. As EMIT only established comprehensive spatial coverage in 2023, no November 2022 data were available over our study regions, and thus only February 2023 was used. In regions with data availability, monthly median hyperspectral reflectance was calculated using GEE.
@@
In ARG, AUS, and CHN, the source maps are binary- or three-class, thus no further class selection or merging was required. However, these crop maps do not include the distribution of non-target crop categories. To capture those classes, we incorporated ESA’s WorldCover [48]; specifically, WorldCover (available via the GEE collection ‘ESA/WorldCover/v200’) pixels labeled as “cropland” but not recognized by the above maps were assigned to the “Other” class.
For crop maps originally provided in vector format, namely QSC (AUS), LGP (BEL), RPG (FRA), CROME (GBR), and BRP (NLD), we converted them into raster format at a spatial resolution of 10 meters and reprojected them to the WGS84 coordinate reference system.
@@
Grid-Based Sampling: To further improve label reliability, we applied a 60-meter inward buffer (based on EMIT resolution) to all reference crop maps after importing them into GEE, which helped eliminate small noisy patches. As shown in Fig. 2 (a), erroneous labels distributed within sample areas are removed after applying buffer filtering. We then matched the buffered crop maps with the filtered satellite data (processed separately for S2 and EMIT) in both space and time. The intersection of valid crop labels and available cloud-free pixels defined the final sampling region, ensuring each pixel had both a class label and satellite data. To reduce sample correlation, the study region was divided into $0.1^{\circ}\times 0.1^{\circ}$ cells in geographic coordinates (WGS84), and random samples were drawn within each cell at a spatial resolution of 60 meters (i.e., ensuring a minimum distance of 60 meters between any two sampled points), as shown in Fig. 2 (b). In USA, one sample was drawn per cell, while in other countries with smaller study areas, five samples were drawn per cell. Taking FRA as an example, a $0.1^{\circ}\times 0.1^{\circ}$ cell is approximately 7.7 km $\times$ 11.1 km, and five uniformly random samples with a minimum of 60-meter spacing remain spatially distinct.
@@ 3 Methodology
To address these challenges, we design a lightweight yet effective CNN model named CropNet, tailored for learning robust features from pixel-level data. Furthermore, to improve the invariance of learned features across geographical domains, where both spectral responses and phenological phases may vary, we introduce dedicated data augmentation strategies for multi-spectral time-series, aimed at enhancing robustness to temporal and spectral perturbations.
@@ 3.1 Construction of Temporal Multi-Spectral Data
1D/2D median features: To construct consistent median features from raw S2 observations, we apply a windowed median compositing strategy. For each pixel, we define a time span of interest with a start day-of-year (DOY) $T_{s}$ and an end DOY $T_{e}$, which typically covers the crop growing season. Although data are collected throughout the year, not all periods contribute equally to crop type classification. The selected time span allows the model to focus on the most relevant phenological phases.
This time span is divided into equal-length temporal windows of size $d$, resulting in $t=\lceil(T_{e}-T_{s})/d\rceil$ discrete bins, where $\lceil\cdot\rceil$ denotes the ceiling operator (rounding up). For each bin, if valid (non-cloudy) observations are available, we compute the median reflectance across all dates within the bin. If no valid observations are found due to cloud coverage, the missing value is filled using linear interpolation [21] along the temporal axis. If the time series contains too many gaps to allow reliable interpolation, the sample is discarded.
Each pixel is ultimately represented as a one-dimensional feature vector of length $10t$, where 10 corresponds to the selected S2 bands, and $t$ is the number of time bins.
@@ 3.2 Augmentation of Temporal Multi-Spectral Data
Time Shift simulates phenological shifts by randomly moving the entire time span forward or backward by up to 10 days (see Section 4.5 for details on parameter selection). Given the original DOY range $T_{s},T_{e}]$, a random integer offset $\Delta_{shift}\in[-10,10]$ is sampled and applied to both start and end DOYs, resulting in a shifted range $T_{s}+\Delta_{shift},T_{e}+\Delta_{shift}]$. Median values are then recalculated over this new time span using the same temporal window $d$ as the original, ensuring feature length consistency. This augmentation accounts for planting date differences or climatic variations across regions.
Time Scale adjusts the temporal extent of the sampling period to capture variability in the duration of crop development. The start and end DOYs of the original time span $T_{s},T_{e}]$ are independently perturbed by random integer offsets $\Delta_{scale}^{s},\Delta_{scale}^{e}\in[-30,10]$ (see Section 4.5 for details on parameter selection). The scaled range is $T_{s}^{\prime},T_{e}^{\prime}]$, where $T_{s}^{\prime}=T_{s}+\Delta_{scale}^{s}$ and $T_{e}^{\prime}=T_{e}+\Delta_{scale}^{e}$, which can expand or contract in length by up to 40 days. To maintain a fixed feature dimension, we compute a new temporal window $d^{\prime}=\lceil(T_{e}^{\prime}-T_{s}^{\prime})/t\rceil$, thereby keeping the total number of bins $t$ unchanged. This augmentation mimics differences in growing season length due to crop varieties, management practices or climate.
Magnitude Warping perturbs the shape of each spectral band’s temporal curve by applying a smooth, multiplicative modulation, aiming to mimic real-world deviations such as sensor noise, sub-pixel heterogeneity, or biomass variability (e.g., the same crop may even have different biomass in different fields) [53]. For a given sample and band $b$, the temporal series is $\mathbf{x}_{b}=(x_{b,1},\ldots,x_{b,t})$. Each time point is scaled by a weight $w_{i}$, i.e., $x^{\prime}_{b,i}=w_{i}\,x_{b,i}$, where the weights are obtained from a cubic spline $S(\cdot)$ fitted to $K$ anchor points $(\kappa_{j},a_{j})$. The anchor abscissae are $\kappa_{j}=1+\frac{j-1}{K-1}(t-1)$ and the ordinates are $a_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(1,\sigma^{2})$ for $j=1,\ldots,K$; the weights are obtained by evaluating the spline at the bin indices $w_{i}=S(i)$ for $i=1,\ldots,t$ [54, 55]. Based on empirical analysis (Section 4.5), we set $\sigma$ to 0.2 and $K$ to 5. To avoid over-regularization, each sample undergoes the same warping transformation across all spectral bands (the same spline $S(\cdot)$ is applied to every band). This augmentation helps the model learn more invariant and generalizable representations under varying growing conditions.
@@ 3.3 Classification Model: CropNet
We design a lightweight CNN, named CropNet, specifically for pixel-level classification of crop type. It performs convolutions jointly along the spectral and temporal dimensions to capture dependencies across both domains, as presented in Fig. 6. CropNet consists of four convolutional blocks, each containing two $3\times 3$ convolutional layers with batch normalization and Rectified Linear Unit (ReLU) activations. Down sampling is performed in the first and third blocks via stride-2 convolutions, retaining more temporal resolution while still allowing the model to capture multi-scale spectral-temporal patterns. After the final convolutional stage, global average pooling reduces the spatial dimensions to $1\times 1$, followed by a linear layer for classification. The output is a probability distribution over crop types.
@@
To improve regularization and prevent overfitting on small per-pixel features, we apply spatial dropout [56] after each convolutional block. Unlike standard dropout, which randomly drops individual values, spatial dropout zeros out entire feature channels, forcing the network to learn more robust and distributed representations across the spectral-temporal space. This is particularly useful in our case, as the input features lack contextual spatial structure and are more prone to overfitting due to limited feature diversity per sample.
We evaluated several architectural variants with different depths and channel widths. Our experiments showed that a four-stage, nine-layer configuration achieves the best trade-off between performance and model complexity. It is a minimal design for optimal generalization. Deeper networks showed no significant gains, while shallower ones consistently underperformed.
For one-dimensional data or data structured as flat time-series vectors, such as hyperspectral reflectance or harmonic coefficients, the 2D convolutions in CropNet can be easily replaced with 1D convolutions. In this configuration, the network architecture remains structurally identical, treating the input as a single-channel signal and performing convolutions along the feature axis.
@@ 4.1 Experimental Setup
Class alignment: Since available crop categories differ across countries and the definition of the “Other” class also varies, we apply a label alignment strategy for fair evaluation. For AUS, ARG, and CHN [42, 41, 45], where the crop maps are binary or ternary and the “Other” class may implicitly include target classes (e.g., soybeans and wheat may exist in AUS “Other”), we adopt the following rule: when using these countries as test regions, any predicted class not explicitly defined within their available categories is reassigned to “Other” before computing accuracy. For example, samples predicted as soybean or wheat in AUS are reassigned to “Other” during evaluation. In the FRA → USA experiment, since FRA contains no sugarcane or cotton samples, we remove these two classes from USA when conducting evaluation; in this case, they cannot simply be merged into “Other” because FRA entirely lacks these categories. For all other transfer scenarios, no special category alignment is applied, as the training regions cover all classes present in the test regions and the definition of “Other” remains consistent.
@@
To evaluate the intrinsic discriminative ability of different features and models without the added complexity of domain shifts, we conduct in-region classification experiments where training and testing samples are taken from the same country. Table 4 summarizes the results where models are trained and tested within the same country. For each country, 80% of samples are randomly selected for training and 20% for testing, with experiments repeated ten times. The results show that median features consistently outperform harmonic coefficients, and 2D median features further outperform 1D median features. Data augmentation remains effective even under these in-region settings, improving both model stability and accuracy, despite limited domain shifts between training and testing data.
It is noteworthy that the FRA crop map is based on farmer-reported data and thus highly reliable, while the USA crop map is automatically generated and its label accuracy is inherently imperfect. Despite this, our model achieves near 99% accuracy on the USA data, suggesting that some degree of overfitting to label noise may have occurred. However, the strong generalization observed in USA-to-other transfer experiments suggests that the learned features still generalize well despite noisy labels.